# Regularized Estimation of Loading Matrix in Factor Models for High-Dimensional Time Series

### v2

@@ 3 Estimation
In this section, we first briefly review a standard estimation method for the loading matrix proposed by Lam et al., (2011) in Section 3.1 and then present our algorithm for obtaining a sparse estimate in Section 3.2.
@@ 3.1 The standard estimation method
$\boldsymbol{\Sigma}_{x}(h)=\frac{1}{n-h}\sum_{t=1}^{n-h}{\rm E}({\mathbf{x}}_{t}{\mathbf{x}}_{t+h}^{\top}),\quad{\mathbf{M}}=\sum_{h=1}^{h_{0}}\boldsymbol{\Sigma}_{x}(h)\boldsymbol{\Sigma}_{x}(h)^{\top},$
[…]
@@
Lam et al., (2011) defined the sample version of these matrices as follows
@@
Thus, the loading space ${\cal M}({\mathbf{A}}^{s})$ is estimated by ${{\cal M}(\widehat{{\mathbf{S}}})}$, where $\widehat{{\mathbf{S}}}=\{\widehat{{\mathbf{s}}}_{1},\ldots,\widehat{{\mathbf{s}}}_{r}\}$ and $\widehat{{\mathbf{s}}}_{i}$ is the eigenvector of $\widehat{{\mathbf{M}}}$ corresponding to the $i$-th largest eigenvalue. In other words,
@@
In practice, $r$ is unknown and needs to be estimated. There are plenty of studies on the estimation of the number of factors; see Bai and Ng, (2002); Onatski, (2009); Kapetanios, (2010); Lam and Yao, (2012); Han et al., (2022). However, this paper focuses on estimating factor loadings rather than $r$, and thus we assume that the number of factors is known throughout the paper.
@@ 3.2.1 The optimization problem
To obtain the desired estimate, the column space of this estimate should be close to ${\cal M}(\widehat{{\mathbf{S}}})$ and the number of nonzero elements in this estimate is sufficiently large. Therefore, we formulate an optimization problem that minimizes the distance between our estimate and ${\cal M}(\widehat{{\mathbf{S}}})$, while incorporating a penalty term that promotes sparsity in the estimate.
@@
where the columns of ${\mathbf{H}}_{i}$ are an orthonormal basis of ${\mathcal{M}}({\mathbf{U}}_{i})$ for $i=1,2$ (Chang et al.,, 2015). It is a quantity between 0 and 1. It is 1 if the two spaces are orthogonal and 0 if ${\mathcal{M}}({\mathbf{U}}_{1})={\mathcal{M}}({\mathbf{U}}_{2})$.
As for the penalty term, we use MCP with the following form: $\mathcal{P}_{\gamma}(x;\lambda)=\lambda|x|-\frac{x^{2}}{2\gamma}$ if $|x|\leq\gamma\lambda$, and $\mathcal{P}_{\gamma}(x;\lambda)=\frac{1}{2}\gamma\lambda^{2}$ if $|x|>\gamma\lambda$, where $\gamma$ is fixed at 3 as in different literature (Breheny and Huang,, 2011, 2015) and $\lambda$ is a tuning parameter, which will be selected based on data-driven criteria.
@@
      subject to $\text{subject to }{\mathbf{s}}_{i}=({\mathbf{I}}-\tilde{{\mathbf{S}}}_{i}\tilde{{\mathbf{S}}}_{i}^{\top}){\mathbf{q}}_{i}\text{ and }\|{\mathbf{s}}_{i}\|_{2}=1.$
@@
Varimax rotation is a widely used technique for improving the interpretability of factor models. It seeks a rotation within ${\cal M}(\widehat{{\mathbf{S}}})$ that maximizes the variance of the squared elements of the loading matrix, i.e.
@@
where $s_{ij}$ is the ($i,j$)-th element in ${\mathbf{S}}$. Consequently, many elements in the loading matrix obtained via varimax rotation are close to zero but not exactly zero. Our method offers two advantages: (1) our estimate has zero elements due to the inclusion of a penalty term in the objective function and (2) we allow the column space of the estimate to deviate slightly from ${\cal M}(\widehat{{\mathbf{S}}})$ in exchange for a more sparse structure.
@@
Uematsu and Yamagata, 2022a investigated the estimation of factors with sparse loadings. As mentioned earlier, the loading matrix is not uniquely defined and can rotate in the loading space. Uematsu and Yamagata, 2022a first chose a specific rotation where the covariance matrix of factors is an identity matrix and the columns in the loading matrix are orthogonal, and then imposed the sparsity assumption on the particular loading matrix. In this study, we relax these assumptions. Our method searches for the most sparse loading matrix in the whole loading space, which yields a more sparse estimate and may provide a clearer interpretation for factors. We compare these two approaches in Section 5 and Section 6. The results show that our method outperforms and also works even when columns of the loading matrix are orthogonal.
@@ 3.2.2 The algorithm
In particular, ${\mathbf{G}}=\widehat{{\mathbf{S}}}\widehat{{\mathbf{S}}}^{\top}$, ${\mathbf{B}}={\mathbf{I}}$ for the problem in (23), and ${\mathbf{B}}={\mathbf{I}}-\tilde{{\mathbf{S}}}_{i}\tilde{{\mathbf{S}}}_{i}^{\top}$ for the problem in (24), which satisfies ${\mathbf{B}}{\mathbf{B}}={\mathbf{B}}$. To use the ADMM algorithm (Boyd et al.,, 2011), we rewrite the optimization problem in (25) as
@@
subject to $\text{subject to }{\mathbf{s}}^{\top}{\mathbf{s}}=1,$
where ${\mathbf{v}}$ is a $p$-dimensional vector containing all the Lagrange multipliers and $\rho$ is a fixed penalty parameter. Here, we set it at 1 as in Ma and Huang, (2017) and Wang et al., (2023). Then, we can update ${\mathbf{s}},{\mathbf{q}},{\mathbf{v}}$ iteratively. At the $(l+1)$-th iteration, given the current values of ${\mathbf{s}}^{(l)},{\mathbf{q}}^{(l)}$, and ${\mathbf{v}}^{(l)}$, the updates of ${\mathbf{s}},{\mathbf{q}},{\mathbf{v}}$ are
@@
$-{\mathbf{s}}^{\top}{\mathbf{G}}{\mathbf{B}}{\mathbf{q}}^{(l)}-{\mathbf{s}}^{\top}{\mathbf{B}}{\mathbf{q}}^{(l)}\rho+{\mathbf{s}}^{\top}{\mathbf{v}}^{(l)},$
@@
(33) can be minimized using the gradient algorithm with the MCP penalty. We use the R package ncvreg (Breheny and Huang,, 2011) to obtain the solution for a fixed value of $\lambda$. Note that $L_{1}$ penalty or SCAD can also be used here.
@@
The stopping criterion is $\|{\mathbf{s}}-{\mathbf{B}}{\mathbf{q}}\|_{2}\leq\epsilon$ as in the literature (Ma and Huang,, 2017; Wang et al.,, 2023), where $\epsilon$ is a small positive value. Here we use $\epsilon=10^{-5}$.
@@
We use BIC to select tuning parameters, which is also used in Uematsu and Yamagata, 2022a . BIC is defined as
@@ 4 Theoretical properties
[…]
@@
[…]
@@
As mentioned earlier, there are two ways to separate the noise process and the factor. The first one assumes that the idiosyncratic error has weak serial dependence and weak cross-sectional dependence with $\sum_{i=1}^{p}\sum_{j=1}^{p}|\sigma_{e,t,ij}|\leq Cp$ for any $t=1,\ldots,n$ and $C$ is a positive constant; see Bai and Ng, (2002), Bai, (2003),Bai and Ng, 2006b , Bai and Ng, (2008), Uematsu and Yamagata, 2022a ,Uematsu and Yamagata, 2022b and among others. The second one assumes that the noise process has no serial dependence but the strong cross-sectional dependence is allowed with $|\sigma_{e,t,ij}|<C$ for any $i,j=1,\ldots,p$ and $t=1,\ldots,n$; see Lam et al., (2011); Lam and Yao, (2012); Chang et al., (2015); Wang et al., (2019); Chen et al., (2022). We follow the second assumption in this study and believe that our approach is also adaptable to the first one; exploring this extension is left for future work. For the second assumption, Conditions (C1)–(C3) and Condition (C5) are quite standard (Lam et al.,, 2011; Lam and Yao,, 2012; Chang et al.,, 2015; Liu and Chen,, 2016; Wang et al.,, 2019; Liu and Zhang,, 2022) and used to ensure that the estimated autocovariance matrices converge.
@@
To measure the strength of factors, Lam and Yao, (2012) introduced a strength factor index $\delta$ and assumed $\|{\mathbf{a}}_{i}^{s}\|^{2}\propto p^{1-\delta}$ for $i=1,\dots,r$ where ${\mathbf{a}}^{s}_{i}$ is the $i$-th column of ${\mathbf{A}}^{s}$ and $\delta\in[0,1]$. When $\delta=0$, the factors are strong; when $\delta>0$, the factors are weak. Chang et al., (2015) proposed a similar measure that is $\|{\mathbf{A}}^{s}\|^{2}_{2}\propto\|{\mathbf{A}}\|_{\min}^{2}\propto p^{1-\delta}$. Since the strength of factors is defined through the norm of loadings, it is reasonable to consider only nonzero elements and replace $p$ with $m$ when imposing the sparsity assumption on the loading matrix in (C4). In fact, $\delta$ reflects the scale of the elements in the loading matrix, and this can be confirmed by the settings of numerical experiments in Lam et al., (2011), where each element in the loading matrix is generated from a uniform random variable on the interval $[-1,1]$ divided by $p^{\delta/2}$. In our paper, under our settings, we will generate the elements in ${\mathbf{A}}^{s}$ independently by standard normal random variables divided by $(m/r)^{\delta/2}$ in the simulation section. We also need to impose an elementwise bound on ${\mathbf{A}}^{s}$ with $\|{\mathbf{A}}^{s}\|_{\max}\leq C_{1}$. In addition, we assume that the sparsity level of loading vectors remain the same. The relationship between $m$ and $p$ definitely plays an important role in the convergence rate of our estimators. We will discuss it in Section 4.
@@
Condition (C6) is a commonly used assumption in models for high-dimensional data analysis, such as regression models in Ma and Huang, (2017) and Wang et al., (2023), and factor models for functional time series, as in Guo et al., (2021) and Fang et al., (2022). Since the tails of a sub-Gaussian random variable are dominated by the tails of a normal random variable, Condition (C6) helps bound the tails of the noise and factor processes. The definition is provided in the Supplemental Materials Section E.
@@
The bounded coherence assumption in Condition (C8) is widely used in matrix theory; see examples in Fan et al., (2018) and Cape et al., (2019). Cape et al., (2019) assumes $\|{\mathbf{S}}\|_{2\rightarrow}\leq C_{\mu}\sqrt{\frac{r}{p}}$ for a $p\times r$ orthonormal matrix ${\mathbf{S}}$. Since ${\mathbf{Q}}$ in our setting is sparse with $m$ nonzero elements, we replace $p$ with $m$ and assume that the sparsity level of ${\mathbf{S}}$ is $O(m)$. Condition (C8) implies that each element in ${\mathbf{S}}$ is bounded by $C_{\mu}\sqrt{\frac{r}{m}}$.
First, we have Theorem 1 for the asymptotic property for $\widehat{{\mathbf{S}}}$ obtained in (21), where is the estimator proposed in Lam et al., (2011) without considering the sparsity.
@@
Theorem 1 shows the impact of the sparsity level on the convergence rate of the estimated loading space by Lam et al., (2011). If the number of nonzero elements in ${\mathbf{Q}}$ grows as fast as the dimension, i.e., $m=O(p)$, ${\cal M}(\widehat{{\mathbf{S}}})$ converges to ${\cal M}({\mathbf{S}})$ at the rate of $p^{\delta}n^{-1/2}$, which is the same as the results in Lam et al., (2011). If $\delta<1$, the more sparse ${\mathbf{Q}}$ is, the more bias $\widehat{{\mathbf{S}}}$ includes, and thus, the slower ${\cal M}(\widehat{{\mathbf{S}}})$ converges to ${\cal M}({\mathbf{S}})$.
Next, we will study the asymptotic properties of our proposed estimator $\widehat{{\mathbf{Q}}}$. The following theorem shows that our estimator converges faster than the one proposed by Lam et al., (2011) when the loading matrix is sparse.
Let $b=\min_{i}\min_{j\in\mathcal{V}_{1}}|q_{ij}|$, which is the minimal signal of ${\mathbf{Q}}$, and $\phi_{n,p,m}=\max\left(m^{2\delta-2}p^{2}n^{-1/2},m^{\delta}\right)$. We define $\tau_{n,p,m}=\phi_{n,p,m}\sqrt{\frac{\log p}{n}}$ if $m=o(p)$, and $\tau_{n,p,m}=p^{\delta}n^{-1/2}$ if $m=O(p)$. We have the following result for the proposed estimator.
@@
Assume that $m_{1}<m_{2}<\ldots<m_{r}$ and $b>a\lambda$ for some constant $a>0$. If $\lambda\gtrsim\tau_{n,p,m}$ and $\tau_{n,p,m}=o(1)$ as $n\rightarrow\infty$ and $p\rightarrow\infty$. Under Conditions (C1)-(C8), then
$\|\widehat{{\bf Q}}-{\bf Q}\|_{2}=O(\tau_{n,p,m})=\begin{cases}O_{p}\left(\phi_{n,p,m}\sqrt{\frac{\log p}{n}}\right)&\text{if }m=o(p),\\ O_{p}\left(m^{\delta-1}pn^{-1/2}\right)=O_{p}\left(p^{\delta}n^{-1/2}\right)&\text{if }m=O(p),\\ \end{cases}$
@@
Theorem 2 shows that the proposed estimators for the loading matrix and the nonzero indexes of the loading matrix are both consistent under some mild conditions. It also demonstrates the impact of the sparsity level of the loading matrix on its estimation error. If $m=O(p)$, $\widehat{{\mathbf{Q}}}$ converges at the same rate as the estimator proposed in Lam and Yao, (2012). If $m=o(p)$, the convergence rate of ${\mathbf{Q}}$ is determined by two terms: the first term, $O(m^{2\delta-2}p^{2}n^{-1/2}\sqrt{\frac{\log p}{n}})$, is related to the estimation bias and the second term, $O(m^{\delta}\sqrt{\frac{\log p}{n}})$, is related to the estimation variance. If $m^{\delta-2}p^{2}n^{-1/2}\geq O(1)$, i.e., the loading matrix is quite sparse, $\widehat{{\mathbf{S}}}$ has a larger bias as shown in Theorem 1, so the estimation error of ${\mathbf{Q}}$ is dominated by the first term. In that case, the more sparse the matrix, the larger the error is. If $m^{\delta-2}p^{2}n^{-1/2}=o(1)$, i.e., the loading matrix is relatively dense, the estimation error is dominated by the second term (variance). In that case, the more sparse the matrix, the smaller the error is.
@@
Theorem 2 shows the estimators are consistent when the sparsity levels of $\{{\mathbf{q}}_{i}\mid i=1,\ldots,r\}$ are all distinct. However, even if it is not true and ${\mathbf{Q}}$ is not uniquely defined, our method can estimate the loading space well and capture one of the most sparse loading matrices; see details in Section 5.
@@
as $n$ and $p$ go to infinity.
Theorem 3 specifies the convergence rate for the estimated common component. If $m=O(p)$, the ratio of the convergence rate of $\widehat{{\bf Q}}\widehat{{\bf z}}_{t}$ to that of $\widehat{{\mathbf{Q}}}$ is the same as that shown in Lam et al., (2011). If $\delta<1$ and $p$ is fixed, the ratio of the convergence rates of these two estimators increases as $m$ increases, which implies that it is more difficult to recover the common component when the loading matrix is quite sparse.

### v3

@@ 3 Estimation
In this section, we first briefly review a standard estimation method for the loading matrix proposed by Lam et al. (2011) in Section 3.1 and then present our algorithm for obtaining a sparse estimate in Section 3.2.
@@ 3.1 The standard estimation method
$\boldsymbol{\Sigma}_{x}(h)=\frac{1}{n-h}\sum_{t=1}^{n-h}{\rm E}({\mathbf{x}}_{t}{\mathbf{x}}_{t+h}^{\top}),\quad{\mathbf{M}}=\sum_{h=1}^{h_{0}}\boldsymbol{\Sigma}_{x}(h)\boldsymbol{\Sigma}_{x}(h)^{\top},$ (20)
where $h_{0}$ is a pre-specified positive integer. Since $\{\mbox{$\varepsilon$}_{t}\}$ has no serial dependence, we have
@@
Lam et al. (2011) defines the sample version of these matrices as follows
@@
Thus, the loading space ${\cal M}({\mathbf{A}}^{s})$ or ${\cal M}({\mathbf{Q}})$ is estimated by ${{\cal M}(\widehat{{\mathbf{S}}})}$, where $\widehat{{\mathbf{S}}}=\{\widehat{{\mathbf{s}}}_{1},\ldots,\widehat{{\mathbf{s}}}_{r}\}$ and $\widehat{{\mathbf{s}}}_{i}$ is the eigenvector of $\widehat{{\mathbf{M}}}$ corresponding to the $i$-th largest eigenvalue. In other words,
@@
In practice, $r$ is unknown and needs to be estimated. There are plenty of studies on the estimation of the number of factors; see Bai and Ng (2002); Onatski (2009); Kapetanios (2010); Lam and Yao (2012). However, the presence of sparsity in the loading matrix may pose a new challenge. Trapani (2018) proposed a randomized sequential procedure to determine the number of factors, which is very general and does not require any assumptions on factors, loadings or errors. Under our settings, the randomized sequential test can be used to select the number of factors.
@@ 3.2.1 The optimization problem
[…]
@@
where the columns of ${\mathbf{H}}_{i}$ are an orthonormal basis of ${\mathcal{M}}({\mathbf{U}}_{i})$ for $i=1,2$ (Chang et al., 2015). It is a quantity between 0 and 1. It is 1 if the two spaces are orthogonal and 0 if ${\mathcal{M}}({\mathbf{U}}_{1})={\mathcal{M}}({\mathbf{U}}_{2})$.
As for the penalty term, we use MCP with the following form: $\mathcal{P}_{\gamma}(x;\lambda)=\lambda|x|-\frac{x^{2}}{2\gamma}$ if $|x|\leq\gamma\lambda$, and $\mathcal{P}_{\gamma}(x;\lambda)=\frac{1}{2}\gamma\lambda^{2}$ if $|x|>\gamma\lambda$, where $\gamma$ is fixed at 3 as in different literature (Breheny and Huang, 2011, 2015) and $\lambda$ is a tuning parameter, which will be selected based on data-driven criteria. Compared with $L_{1}$-norm penalty term used in Wu et al. (2026), MCP satisfies the oracle property. MCP provides asymptotically unbiased estimation while $L_{1}$-norm penalty term cannot.
@@
      $\text{subject to }{\mathbf{s}}_{i}=({\mathbf{I}}-\tilde{{\mathbf{S}}}_{i}\tilde{{\mathbf{S}}}_{i}^{\top}){\mathbf{q}}_{i}\text{ and }\|{\mathbf{s}}_{i}\|_{2}=1.$
@@
[…]
@@
[…]
@@
In the literature, there are several studies on factor models with sparse loadings (Kristensen, 2017; Pelger and Xiong, 2022; Uematsu and Yamagata, 2022b, a; Wu et al., 2026). As mentioned earlier, the loading matrix is not uniquely defined and can rotate in the loading space. All existing work first chooses a specific rotation where the covariance matrix of factors is an identity matrix and/or the columns in the loading matrix are orthogonal, and then impose the sparsity assumption on the particular loading matrix. In this study, we relax these assumptions. Our method searches for the most sparse loading matrix in the whole loading space, which yields a more sparse estimate for a more parsimonious model to improve model interpretation and predictive performance. We compare these two approaches in Section 5 and Section 6. The results show that our method outperforms others.
@@ 3.2.2 The algorithm
In particular, $^{\mathbf{G}}=\widehat{{\mathbf{S}}}\widehat{{\mathbf{S}}}^{\top}$, ${\mathbf{B}}={\mathbf{I}}$ for the problem in (24), and ${\mathbf{B}}={\mathbf{I}}-\tilde{{\mathbf{S}}}_{i}\tilde{{\mathbf{S}}}_{i}^{\top}$ for the problem in (25), which satisfies ${\mathbf{B}}{\mathbf{B}}={\mathbf{B}}$. To use the ADMM algorithm (Boyd et al., 2011), we rewrite the optimization problem in (26) as
@@
$\text{subject to }{\mathbf{s}}^{\top}{\mathbf{s}}=1,$
where $\mathbf{v}$} is a $p$-dimensional vector containing all the Lagrange multipliers and $\rho$ is a fixed penalty parameter. Here, we set it at 1 as in Ma and Huang (2017) and Wang et al. (2023). Then, we can update ${\mathbf{s}},{\mathbf{q}},{\mathbf{v}}$ iteratively. At the $(l+1)$-th iteration, given the current values of ${\mathbf{s}}^{(l)},{\mathbf{q}}^{(l)}$, and ${\mathbf{v}}^{(l)}$, the updates of ${\mathbf{s}},{\mathbf{q}},{\mathbf{v}}$ are
@@
${\mathbf{s}}^{\top}{\mathbf{G}}{\mathbf{B}}{\mathbf{q}}^{(l)}-{\mathbf{s}}^{\top}{\mathbf{B}}{\mathbf{q}}^{(l)}\rho+{\mathbf{s}}^{\top}{\mathbf{v}}^{(l)},$
@@
(34) can be minimized using the gradient algorithm with the MCP penalty. We use the R package ncvreg (Breheny and Huang, 2011) to obtain the solution for a fixed value of $\lambda$.
@@
The stopping criterion is $\|{\mathbf{s}}-{\mathbf{B}}{\mathbf{q}}\|_{2}\leq\epsilon$ as in the literature (Ma and Huang, 2017; Wang et al., 2023), where $\epsilon$ is a small positive value. Here we use $\epsilon=10^{-5}$.
@@
We use BIC to select tuning parameters, which is also used in Uematsu and Yamagata (2022b). BIC is defined as
@@ 4 Theoretical properties
    $\mbox{$\varepsilon$}_{t}$ and ${\mathbf{f}}_{t}^{s}$ are uncorrelated given ${\cal F}_{-\infty}^{t-1}$. Let $\boldsymbol{\Sigma}_{e,t}$ be the covariance of $\mbox{$\varepsilon$}_{t}$. $|\sigma_{e,t,ij}|<\boldsymbol{\Sigma}_{\epsilon}^{2}<\infty$ for $i,j=1,\ldots,p$, and $t=1,\ldots,n$. In other words, the absolute value of each element of $\boldsymbol{\Sigma}_{e,t}$ remains bounded by a constant $\sigma_{\epsilon}^{2}$ as $p$ increases to infinity, for $t=1,\ldots,n$.
@@
    $\mbox{$\varepsilon$}_{t}$’s are independent sub-Gaussian random vectors. Each random vector in the sequence $\{{\mathbf{f}}_{t}^{s}\}$ follows a sub-Gaussian distribution.
@@
As mentioned earlier, there are two ways to separate the noise process and the factor. The first one assumes that the idiosyncratic error has weak serial dependence and weak cross-sectional dependence with $\sum_{i=1}^{p}\sum_{j=1}^{p}|\sigma_{e,t,ij}|\leq Cp$ for any $t=1,\ldots,n$ and $C$ is a positive constant; see Bai and Ng (2002), Bai (2003), Bai and Ng (2006a), Bai and Ng (2008), Uematsu and Yamagata (2022b),Uematsu and Yamagata (2022a) and among others. The second one assumes that the noise process has no serial dependence but the strong cross-sectional dependence is allowed with $|\sigma_{e,t,ij}|<C$ for any $i,j=1,\ldots,p$ and $t=1,\ldots,n$; see Lam et al. (2011); Lam and Yao (2012); Chang et al. (2015); Wang et al. (2019); Chen et al. (2022). We follow the second assumption in this study and believe that our approach is also adaptable to the first one; exploring this extension is left for future work. For the second assumption, Conditions (C1)–(C3) and Condition (C5) are quite standard (Lam et al., 2011; Lam and Yao, 2012; Chang et al., 2015; Liu and Chen, 2016; Wang et al., 2019; Liu and Zhang, 2022) and used to ensure that the estimated autocovariance matrices converge.
@@
To measure the strength of factors, Lam and Yao (2012) introduced a strength factor index $\delta$ and assumed $\|{\mathbf{a}}_{i}^{s}\|^{2}\propto p^{1-\delta}$ for $i=1,\dots,r$, where ${\mathbf{a}}^{s}_{i}$ is the $i$-th column of ${\mathbf{A}}^{s}$ and $\delta\in[0,1]$. When $\delta=0$, the factors are strong; when $\delta>0$, the factors are weak. Chang et al. (2015) proposed a similar measure that is $\|{\mathbf{A}}^{s}\|^{2}_{2}\propto\|{\mathbf{A}}\|_{\min}^{2}\propto p^{1-\delta}$. Since the strength of factors is defined through the norm of loadings, it is reasonable to consider only nonzero elements and replace $p$ with $m$ when imposing the sparsity assumption on the loading matrix in (C4). In fact, $\delta$ reflects the scale of the elements in the loading matrix, and this can be confirmed by the settings of numerical experiments in Lam et al. (2011), where each element in the loading matrix is generated from a uniform random variable on the interval [1,1][-1,1] divided by $p^{\delta/2}$. In simulation section, we will generate random variables from a truncated standard normal distribution and then divide them by $(m/r)^{\delta/2}$ to obtain elements in ${\mathbf{A}}^{s}$. In addition, we impose an elementwise bound on ${\mathbf{A}}^{s}$ with $\|{\mathbf{A}}^{s}\|_{\max}\leq C_{1}$ and assume that the sparsity level of loading vectors remains the same in Condition (C4). The relationship between $m$ and $p$ definitely plays an important role in the convergence rate of our estimators. We will discuss it in Section 4.
@@
Condition (C6) is a commonly used assumption in models for high-dimensional data analysis, such as regression models in Ma and Huang (2017) and Wang et al. (2023), and factor models for functional time series, as in Guo et al. (2021) and Fang et al. (2022). Since the tails of a sub-Gaussian random variable are dominated by the tails of a normal random variable, Condition (C6) helps bound the tails of the noise and factor processes. The definition is provided in the Supplemental Materials Section E.
@@
The bounded coherence assumption in Condition (C8) is widely used in matrix theory; see examples in Fan et al. (2018) and Cape et al. (2019). Cape et al. (2019) assumes $\|{\mathbf{S}}\|_{2\rightarrow}\leq C_{\mu}\sqrt{\frac{r}{p}}$ for a $p\times r$ orthonormal matrix $\mathbf{S}$}. Since $\mathbf{Q}$} in our setting is sparse with $m$ nonzero elements, we replace $p$ with $m$ and assume that the sparsity level of $\mathbf{S}$} is $O(m)$. Condition (C8) implies that each element in $\mathbf{S}$} is bounded by $C_{\mu}\sqrt{\frac{r}{m}}$.
First, we have Theorem 1 for the asymptotic property for $\widehat{{\mathbf{S}}}$ obtained in (22), where is the estimator proposed in Lam et al. (2011) without considering the sparsity.
@@
[…]
Next, we will study the asymptotic properties of our proposed estimator $\widehat{{\mathbf{Q}}}$. The following theorem shows that our estimator converges faster than the one proposed by Lam et al. (2011) when the loading matrix is sparse.
Let $b=\min_{i}\min_{j\in\mathcal{V}_{i}}|q_{ij}|$, which is the minimal signal of $\mathbf{Q}$}, and $\phi_{n,p,m}=\max\left(m^{2\delta-2}p^{2}n^{-1/2},m^{\delta}\right)$. We define $\tau_{n,p,m}=\phi_{n,p,m}\sqrt{\frac{\log p}{n}}$ if $m=o(p)$, and $\tau_{n,p,m}=p^{\delta}n^{-1/2}$ if $m=O(p)$. We have the following result for the proposed estimator.
@@
Assume that $m_{1}<m_{2}<\ldots<m_{r}$ and $b>\gamma\lambda$. If $\lambda\gtrsim\tau_{n,p,m}$ and $\tau_{n,p,m}=o(1)$ as $n\rightarrow\infty$ and $p\rightarrow\infty$. Under Conditions (C1)-(C8), then
[…]
@@
Theorem 2 shows that the proposed estimators for the loading matrix and the nonzero indexes of the loading matrix are both consistent under some mild conditions. It also demonstrates the impact of the sparsity level of the loading matrix on its estimation error. If $m=O(p)$, $\widehat{{\mathbf{Q}}}$ converges at the same rate as the estimator proposed in Lam et al. (2011). If $m=o(p)$, the convergence rate of $\widehat{{\mathbf{Q}}}$ is determined by two terms: the first term, $O(m^{2\delta-2}p^{2}n^{-1/2}\sqrt{\frac{\log p}{n}})$, is related to the estimation bias and the second term, $O(m^{\delta}\sqrt{\frac{\log p}{n}})$, is related to the estimation variance. If $m^{\delta-2}p^{2}n^{-1/2}\geq O(1)$, i.e., the loading matrix is quite sparse, $\widehat{{\mathbf{S}}}$ has a larger bias as shown in Theorem 1, so the estimation error of $\mathbf{Q}$} is dominated by the first term. In that case, the more sparse the matrix, the larger the error is. If $m^{\delta-2}p^{2}n^{-1/2}=o(1)$, i.e., the loading matrix is relatively dense, the estimation error is dominated by the second term (variance). In that case, the more sparse the matrix, the smaller the error is.
@@
Theorem 2 shows the estimators are consistent when the sparsity levels of $\{{\mathbf{q}}_{i}\mid i=1,\ldots,r\}$ are all distinct. However, even if it is not true and $\mathbf{Q}$} is not uniquely defined, our method can estimate the loading space well and capture one of the most sparse loading matrices; see the example in Section 5.1.
@@
as $n$ and $p$ go to infinity, where $\widehat{{\bf z}}_{t}=(\widehat{{\mathbf{Q}}}^{\top}\widehat{{\mathbf{Q}}})^{-1}\widehat{{\mathbf{Q}}}^{\top}{\mathbf{x}}_{t}$.
Theorem 3 specifies the convergence rate for the estimated common component. If $m=O(p)$, the rate is the same as that shown in Lam et al. (2011), and the right-hand side of (36) is $O_{p}(p^{-\delta/2}\|\widehat{{\mathbf{Q}}}-{\mathbf{Q}}\|_{2}+p^{-1/2})$. If $m=o(p)$, the right-hand side of (36) can be written as $O_{p}(\sqrt{\frac{m}{p}}m^{-\delta/2}\|\widehat{{\mathbf{Q}}}-{\mathbf{Q}}\|_{2}+p^{-1/2})$, which is smaller than the one in Lam et al. (2011), and our method produces a better estimator than Lam et al. (2011).